OpenStudy (anonymous):

how do you rewrite this exspression whithout using grouping symbols -6(8+12)

OpenStudy (anonymous):

I'm not quite sure, but wouldn't it be -6+-120?

OpenStudy (anonymous):

Parentheses ( ): Parentheses are the most commonly used symbols for grouping. So you want to get rid of those. This is done by multiplying -6 into 8 and 12. \[-6(8+12)=8*(-6)+12*(-6)=-48-72=-120\]

OpenStudy (anonymous):

Other grouping symbols: Brackets [ ] and Braces { }: Brackets and braces are also used frequently for grouping and have the same effect as parentheses. Using the different types of symbols helps when there’s more than one grouping in a problem. It’s easier to tell where a group starts and ends. Radical √: Also called the square root symbol, this is used for finding roots. Fraction Line (Vinculum) —: The fraction line also acts as a grouping symbol. Everything above the line in the numerator is grouped together, and everything below the line in the denominator is grouped together. - Source: dummies.com